OpenStudy (anonymous):

What are the coordinates of the vertex of y = x^2 - 4x - 5? (2, -9) (2, -5) (-2, 7) (-2, -5)

OpenStudy (anonymous):

y=−2x−5

OpenStudy (anonymous):

-2, -5 :) Please give a medal

OpenStudy (anonymous):

Or fan if this helped

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem. :D

OpenStudy (anonymous):

i think it is a good thing to show your work and also don't you need to complete the square then turn into vertex form

OpenStudy (anonymous):

What? @helpme1.2 ???

OpenStudy (anonymous):

vertex form: y = a(x-h)^2+k y = x^2 - 4x - 5 y = (x^2-4x +4)-5-4 y = (x-2)^2-9 now that is in vertex form which means the coordinates would be (2,9)

OpenStudy (anonymous):

(2 -9)*

OpenStudy (anonymous):

does this make sense @Nikkeyiboo ?